Hello and thanks so much for your time.

Write the equation of the line passing through each of the given points with the indicated slope. Give your results in slope-intercept form, where possible. A note by this problem says to rewrite it in standard form.
(0,5),m= 3/5
McGraw Hill 6th edition, 7.3, Problem #6. Thanks!!

Answer by stanbon(75887)

Give your results in slope-intercept form, where possible. A note by this problem says to rewrite it in standard form.
(0,5),m= 3/5
-----------
You are given the slope and the y-intercept
EQUATION:
y=(3/5)x+5
This is the slope-intercept form.
To get the standard form you need ax+by=c
Multiply thru by 5 to get 5y=3x+25
Re-arrange the terms as follows:
3x-5y=-25
Cheers,
Stan H.